The following labor standards have been established for a particular product:

Standard labor-hours per unit of output 10.1 hours
Standard labor rate $ 13.90 per hour

The following data pertain to operations concerning the product for the last month:

Actual hours worked 7,900 hours
Actual total labor cost $ 106,650
Actual output 1,100 units

What is the labor efficiency variance for the month?

Multiple Choice

  • $47,779 F

  • $47,779 U

  • $43,335 F

  • $44,619 F
Answer #1

Labor efficiency variance

= (SH-AH) *SR

= (1100*10.1 - 7900)*13.90

= (11,110 - 7900)*13.90

= 44,619 F